# Planning with Adaptive World Models for Autonomous Driving

### v1

@@ 4 Experiments
In this section, we compare several baselines with AdaptiveDriver on the nuPlan benchmark and show that our proposed approach achieves state-of-the-art closed-loop planning performance. Further, we demonstrate that AdaptiveDriver generalizes to never-before-seen cities.
@@
nuPlan Dataset. We evaluate all methods on the nuPlan benchmark [3], which includes $\sim$10M driving logs, collected in Boston, Pittsburgh, Las Vegas and Singapore. nuPlan identifies $73$ types of interesting scenarios e.g. {changing lanes, starting left turn, unprotected right turn, etc.}, but only evaluates $14$ scenarios in the official benchmark.
@@
Evaluation Setup and Metrics. The nuPlan benchmark evaluates planners under three evaluation settings: C1, C2 and C3. Following prior work [34], we focus on C2 and C3 as these metrics directly evaluate planning-centric performance. Specifically, C2 and C3 measure ego progress along expert trajectory, speed limit compliance, driving direction compliance, time to collision within bounds, and ego is comfortable. The overall score is computed using a weighted sum of the above metrics. We refer the reader to  [3] for further details.
       Model        City        C2        C3        PDM-C [4]        PIT        92.94        90.59        BOS        95.56        94.45        SIN        93.49        94.15        LAS        95.95        95.21        AdaptiveDriver w/ City-BehaviorNet (Ours)        PIT        91.48        91.94        BOS        96.32        96.21        SIN        95.31        95.29        LAS        96.75        97.08        AdaptiveDriver w/ Log-BehaviorNet (Ours)        PIT        92.39        92.42        BOS        96.39        96.44        SIN        95.22        95.30        LAS        96.61        97.25
@@
Table 3: Per-City Breakdown Analysis. We observe that all planners with reactive world models (e.g. AdaptiveDriver w/ City-BehaviorNet and AdaptiveDriver w/ Log-BehaviorNet) consistently outperform PDM-C in closed-loop (e.g. reactive) planning performance (C3). However, reactive world models do not consistently improve open-loop (e.g. non-reactive) planning performance (C2).
@@
Comparison to State-of-the-Art. We evaluate several rule-based and learning-based planners on the nuPlan test set. First, we confirm Dauner et. al.’s [4] observation that rule-based methods like IDM and PDM-C outperform prior learning-based methods like Raster Model, LaneGCN, and UrbanDriver. Next, we note that both variants of AdaptiveDriver (w/ City-BehaviorNet and w/ Log-BehaviorNet) beat PDM-C on C2 and C3. Notably, AdaptiveDriver w/ City-BehaviorNet improves over PDM-C by $0.48\%$ on C2 and $1.52\%$ on C3 compared (cf. Table 2). Similarly, AdaptiveDriver w/ Log-BehaviorNet showcases an improvement of $0.6\%$ on C2 and $1.71\%$ on C3 over the PDM-C baseline. Lastly, we present a per-city breakdown analysis in Table 3. We find that all planners with reactive world models consistently outperform the baseline in closed-loop planning performance (C3), with improvements in performance of up to $1.74\%$. Further, we confirm the (obvious fact) that motion planning for AVs is more difficult in some cities than in others. For example, although we achieve more than 97% closed-loop planning accuracy in LAS, we only achieve 92% performance in PIT. However, we note that planning performance is largely saturated on the nuPlan benchmark, suggesting that future work should explore more challenging scenario types.
       Model        City        C2        C3        AdaptiveDriver w/ City-Oracle        All        95.10        95.20        AdaptiveDriver w/ City-BehaviorNet        All        94.97 (-0.13)        95.13 (-0.07)        AdaptiveDriver w/ Log-Oracle        All        95.29        95.47        AdaptiveDriver w/ Log-BehaviorNet        All        95.15 (-0.14)        95.35 (-0.12)
@@
Table 4: Oracle Target IDM Parameters. We evaluate the quality of BehaviorNet’s predictions compared to oracle target IDM parameters. Specifically, City-Oracle and Log-Oracle always use the target IDM parameters (rather than predicted IDM parameters) that minimizes Eq. 1 for each log in world model rollouts. Notably, BehaviorNet closely matches the performance of the oracle on the test set, suggesting that our model accurately estimates IDM control parameters from scene context.
       Train Cities        Test City        C2        C3        PIT, BOS, SIN        LAS        96.55 (-0.06)        97.21 (-0.04)        All        LAS        96.61        97.25        PIT, BOS, LAS        SIN        95.05 (-0.17)        95.12 (-0.08)        All        SIN        95.22        95.20        BOS, SIN, LAS        PIT        91.60 (-0.79)        92.02 (-0.40)        All        PIT        92.39        92.42        PIT, SIN, LAS        BOS        96.22 (-0.17)        96.23 (-0.21)        All        BOS        96.39        96.44
@@
Table 5: AdaptiveDriver Generalizes to New Cities. Although learning-based planners can leverage data-driven priors, such methods may fail when evaluated in out-of-distribution scenarios. We evaluate AdaptiveDriver’s ability to generalize to previously unseen cities by holding out city-specific sequences when training BehaviorNet. We find that the planning performances (C2 and C3) are competitive with planners trained on all cities, indicating AdaptiveDriver’s generalizability.
@@
Figure 5: Qualitative Comparison of AdaptiveDriver vs. PDM-C. We evaluate the performance of our AdaptiveDriver against the current state-of-the-art in the simulation of two test scenarios (left and right) from the nuPlan dataset. These scenarios are 15 seconds long and sampled at 10 Hz. Timestamps are displayed on both the top and bottom rows to illustrate collision events. The yellow box represents the ego vehicle and the gray boxes represent other agents. According to nuPlan’s simulation logic, vehicles that collide with each other are removed from the simulation. On the left, the ego vehicle moves from a parked position onto the road. AdaptiveDriver initially slows down to avoid a collision before smoothly entering the road, while PDM-C collides. On the right, we observe an agent about to cross the road. PDM-C drives slowly and collides, whereas our planner safely navigates forward. Please see the supplement for videos.
@@
Oracle Target Parameters vs. BehaviorNet Predicted Parameters. We assess the quality of BehaviorNet’s predictions by comparing them to target IDM parameters from Eq. 1. The planning performance of AdaptiveDriver w/ Log-BehaviorNet closely approximates that of BehaviorNet w/ Log-Oracle, with differences of $0.14$ and $0.12$ in C2 and C3, respectively. The minimal performance difference between AdaptiveDriver using City-Oracle vs. City-BehaviorNet, and Log-Oracle vs. Log-BehaviorNet in Table 4, suggests that our approach can reliably predict future agent actions.
@@
AdaptiveDriver Generalizes to New Cities. As illustrated in Figure 3 (b), we observe that many agent behaviors are shared across cities. For instance, aggressive drivers in Boston and Pittsburgh may exhibit similar tailgating behaviors. To validate this observation, we train AdaptiveDriver w/ Log-BehaviorNet on three cities and test it on a held-out city (cf. Table 5). We find that the planning scores on C2 and C3 for seen-vs-unseen cities are comparable, suggesting that AdaptiveDriver can generalize beyond city boundaries.
5 Conclusion
@@
6 Acknowledgements
We would like to thank Brian Yang for his feedback on drafts. This work was supported in part by funding from Bosch Research and the NSF GRFP (Grant No. DGE2140739).
Appendix A Implementation Details
We build upon the codebases of PDM-C [4] and nuPlan [3] for this work. We modify PDM-C [4] by adding additional longitudinal velocities and lateral offsets to the reference path. In addition, we generate additional trajectory proposals by modulating min gap to lead agents, headway time, maximum acceleration and maximum deceleration. In total, we generate $150$ proposals per timestep. We train BehaviorNet for $10$ epochs using the Adam optimizer [35] with a learning rate of $5\mathrm{e}{-5}$. We select a map context of radius $R=100$m. We present BehaviorNet’s architecture in Figure 6.
Appendix B Visualizing the Quality of Log-Specific Target IDM Parameters
We visualize the min-gap distribution of Log-BehaviorNet’s rollouts in Figure 7 and find that our model more closely matches the distribution of real driving logs compared to IDM.
Figure 7: Log-BehaviorNet Rollouts Closely Resemble Real-World Distributions. We compare the gap between the ego-vehicle and lead agent (min-gap) in PIT (Pittsburgh) and SIN (Singapore) across a) real-world logs, b) Default IDM (a reactive world model optimized over the entire dataset), and c) our proposed scenario-specific world model used in AdaptiveDriver. We observe that our model’s min-gap distribution closely aligns with the real-world distribution, outperforming prior methods, as shown in Table 2.
Appendix C AdaptiveDriver-Hybrid: Combining Open-Loop Forecasting with Closed-Loop Planning
We evaluate several rule-based and learning-based planners on the nuPlan test set in Table 2 of the main paper. We extend these results by also evaluating on the C1 metric in Table 6. First, we note that rule-based planners like IDM and PDM-C outperform prior learning-based methods like Raster Model, LaneGCN, and UrbanDriver on C2 and C3. In contrast, learning-based planners like PDM-O and UrbanDriver outperforms rule-based planners on C1. We observe the same trend with our proposed models: AdaptiveDriver w/ City-BehaviorNet and AdaptiveDriver w/ Log-BehaviorNet achieve C1 scores of $65.42$ and $64.29$ respectively, while PDM-O and UrbanDriver achieve C1 scores of $80.01$ and $73.37$ respectively.
Similar to PDM-H [4], we construct a hybrid model that combines AdaptiveDriver’s predictions for short horizon planning (first two seconds) and PDM-O’s predictions for long-term waypoints (greater than two seconds). This hybrid model (AdaptiveDriver-Hybrid w/ Log-BehaviorNet) outperforms all other methods on C2 and C3 metrics and achieves competitive performance on C1 metric.
       Model        City        C1        C2        C3        Raster Model  [32]        All        64.76        78.50        75.67        LaneGCN [33]        All        69.44        78.52        77.12        UrbanDriver [7]        All        73.37        80.79        81.16        IDM [5]        All        60.56        83.57        81.61        PDM-O [4]        All        80.01        77.70        79.36        PDM-C [4]        All        67.23        94.49        93.61        AdaptiveDriver w/ City-BehaviorNet (Ours)        All        65.42        94.97        95.13        AdaptiveDriver w/ Log-BehaviorNet (Ours)        All        64.29        95.15        95.35        AdaptiveDriver-Hybrid w/ Log-BehaviorNet (Ours)        All        78.84        95.15        95.35
Table 6: Additional nuPlan Benchmark Results. We extend benchmarking from Table 2 of the main paper by adding C1 metric results for completeness. We evaluate rule-based and learning-based planners on the nuPlan test set. In contrast to Table 2, where rule-based planners like IDM and PDM-C outperform prior learning-based methods like Raster Model, LaneGCN, and UrbanDriver on C2 and C3, learning based planners outperforms rule-based methods on C1. Following prior work [4], we build a hybrid model that combines AdaptiveDriver with PDM-O, outperforming all prior work on C2 and C3, with significantly improved performance (compared to AdaptiveDriver) on C1. We refer the reader to nuPlan  [3] for metric definitions.
Appendix D BehaviorNet Regression Baseline
In the main paper, we train BehaviorNet with a K-way softmax loss to predict agent behaviors as IDM parameters from one of $K$ behavioral clusters. However, instead of discretizing BehaviorNet’s predictions into one of $K$ classes, one can train a regressor to directly predict IDM control parameters. We ablate the training of BehaviorNet as a regressor in Table 7. Notably, the performance of AdaptiveDriver w/ Log-BehaviorNet (Regressor) is competitive with AdaptiveDriver w/ Log-BehaviorNet (Classifier). However, we posit that training BehaviorNet to directly predict control parameters may generalize better to larger numbers of behavioral clusters.
       Model        City        C2        C3        AdaptiveDriver w/ Log-BehaviorNet (Classifier)        All        95.15        95.35        AdaptiveDriver w/ Log-BehaviorNet (Regressor)        All        95.20        95.27
Table 7: Regression Baseline. We frame BehaviorNet’s training objective as a K-way classification task to predict agent behavior as IDM controls in the main paper. Instead, one could train a regressor to directly predict control parameters. We note that AdaptiveDriver w/ Log-BehaviorNet (Regressor)’s performance (C3 of $95.27$) is competitive to that of the baseline (C3 of $95.35$).
Appendix E Limitations and Future Work.
Don’t City-Specific Models Add Complexity and Limit Generalization? No! Most autonomy stacks are already city-specific because they make use of city-specific maps. In practice, the overhead of adding additional behavior parameters per city is minimal. More importantly, we demonstrate that AdaptiveDriver captures prototypical agent behaviors (e.g., aggressive vs. passive) that can generalize to never-before-seen cities.
What about Pedestrians and Cyclists? For simplicity, AdaptiveDriver only predicts reactive world model rollouts for cars. Pedestrians and cyclists use constant velocity rollouts (like in PDM-C). Future work should integrate reactive policies for cars, cyclists and pedestrians to better represent multi-agent interactions.
@@ Appendix D BehaviorNet Regression Baseline
When does AdaptiveDriver Fail? Since our world models build on IDM’s PID controller, we inherit its flaws. We note that IDM can be too conservative (when it mistakes a parked vehicle for a lead vehicle and stops) or too aggressive (when traveling at high speeds along a curve. Secondly, IDM only interacts with the lead vehicle, limiting multi-agent reasoning. Although we primarily focus on building city-specific and log-specific world models using the IDM, modeling agent-specific behavior models can potentially yield better performance. Future work should explore alternative optimization functions to maximize distribution similarity between real driving logs and simulated rollouts for learning control parameters.

### v3

@@ IV Experiments
In this section, we compare AdaptiveDriver with several baselines on the nuPlan benchmark and show that our proposed approach achieves state-of-the-art closed-loop planning performance. Further, we demonstrate that AdaptiveDriver generalizes to never-before-seen cities.
@@
nuPlan Dataset. We evaluate all methods on the nuPlan benchmark [3], which includes $\sim$10M driving logs, collected in Boston, Pittsburgh, Las Vegas and Singapore. nuPlan identifies $73$ types of interesting scenarios e.g. {changing lanes, starting left turn, unprotected right turn, etc.}, but only evaluates $14$ scenarios in the official benchmark. We evaluate on the official Val-14 and Test-14 Hard [36] splits in Table II, and evaluate all ablations on a representative mini val-set. We construct the mini val-set by sampling 10 scenarios from all 14 scenario types across all 4 cities (560 logs).
@@
Evaluation Setup and Metrics. The nuPlan benchmark evaluates planners under three evaluation settings: OLS, NR-CLS and R-CLS. Following prior work [36, 40, 39], we focus on NR-CLS and R-CLS as these metrics directly evaluate planning-centric performance. Specifically, NR-CLS and R-CLS measure ego progress along expert trajectory, speed limit compliance, driving direction compliance, time to collision within bounds, and ego is comfortable. The overall score is computed using a weighted sum of the above metrics. We refer the reader to  [3] for further details.
Implementation Details. We build upon the codebases of PDM-C [4] and nuPlan [3] for this work. AdaptiveDriver adds an IDM controller on top of PDM-C’s [4] constant velocity world model rollouts. We train BehaviorNet for $10$ epochs using the Adam optimizer [41] with a learning rate of $5\mathrm{e}{-5}$. We select a map context of radius $R=100$m. Our code is available on GitHub.
       Model        City        NR-CLS        R-CLS        PDM-C [4]        PIT        92.94        90.59        BOS        95.56        94.45        SIN        93.49        94.15        LAS        95.95        95.21        AdaptiveDriver w/ City-BehaviorNet (Ours)        PIT        91.48        91.94        BOS        96.32        96.21        SIN        95.31        95.29        LAS        96.75        97.08        AdaptiveDriver w/ Log-BehaviorNet (Ours)        PIT        92.39        92.42        BOS        96.39        96.44        SIN        95.22        95.30        LAS        96.61        97.25
@@
TABLE III: Per-City Breakdown Analysis. We observe that all planners with reactive world models (e.g. AdaptiveDriver w/ City-BehaviorNet and AdaptiveDriver w/ Log-BehaviorNet) consistently outperform PDM-C in closed-loop (e.g. reactive) planning performance (R-CLS) on the mini val-set. However, reactive world models do not consistently improve open-loop (e.g. non-reactive) planning performance (NR-CLS).
@@
Comparison to State-of-the-Art. We evaluate several rule-based and learning-based planners on the nuPlan test set. First, we find that AdaptiveDriver w/ Log-BehaviorNet showcases 1.7% and 3.7% improvement on R-CLS over the PDM-C baseline on Val-14 and Test-14 Hard respectively. Notably, our approach outperforms more complex LLM-based planers [38], and convincingly outperforms our PDM-C baseline on Test-14 Hard. We present a per-city breakdown analysis in Table III. Interestingly, we find that all planners with reactive world models consistently outperform the baseline in closed-loop planning performance (R-CLS), with improvements in performance of up to $1.74\%$. Further, we confirm the (obvious fact) that motion planning for AVs is more difficult in some cities than in others. For example, although we achieve more than 97% closed-loop planning accuracy in LAS, we only achieve 92% performance in PIT.
       Model        City        NR-CLS        R-CLS        AdaptiveDriver w/ City-Oracle        All        95.10        95.20        AdaptiveDriver w/ City-BehaviorNet        All        94.97 (-0.13)        95.13 (-0.07)        AdaptiveDriver w/ Log-Oracle        All        95.29        95.47        AdaptiveDriver w/ Log-BehaviorNet        All        95.15 (-0.14)        95.35 (-0.12)
@@
TABLE IV: Oracle Target IDM Parameters. We evaluate the quality of BehaviorNet’s predictions compared to oracle target IDM parameters on the mini val-set. Specifically, City-Oracle and Log-Oracle always use the target IDM parameters (rather than predicted IDM parameters) that minimizes Eq. 1 for each log in world model rollouts. Notably, BehaviorNet closely matches the performance of the oracle on the test set, suggesting that our model accurately estimates IDM control parameters from scene context.
       Train Cities        Test City        NR-CLS        R-CLS        PIT, BOS, SIN        LAS        96.55 (-0.06)        97.21 (-0.04)        All        LAS        96.61        97.25        PIT, BOS, LAS        SIN        95.05 (-0.17)        95.12 (-0.08)        All        SIN        95.22        95.20        BOS, SIN, LAS        PIT        91.60 (-0.79)        92.02 (-0.40)        All        PIT        92.39        92.42        PIT, SIN, LAS        BOS        96.22 (-0.17)        96.23 (-0.21)        All        BOS        96.39        96.44
@@
TABLE V: AdaptiveDriver Generalizes to New Cities. Although learning-based planners can leverage data-driven priors, such methods may fail when evaluated in out-of-distribution scenarios. We evaluate AdaptiveDriver’s ability to generalize to previously unseen cities in the mini val-set by holding out city-specific sequences when training BehaviorNet. We find that the planning performances (NR-CLS and R-CLS) are competitive with the model trained on all cities, indicating AdaptiveDriver’s generalizability.
@@
Figure 6: Qualitative Comparison of AdaptiveDriver vs. PDM-C. We evaluate the performance of our AdaptiveDriver against the current state-of-the-art in the simulation of two test scenarios (left and right) from the nuPlan dataset. These scenarios are 15 seconds long and sampled at 10 Hz. Timestamps are displayed on both the top and bottom rows to illustrate collision events. The yellow box represents the ego vehicle and the gray boxes represent other agents. According to nuPlan’s simulation logic, vehicles that collide with each other are removed from the simulation. On the left, the ego vehicle moves from a parked position onto the road. AdaptiveDriver initially slows down to avoid a collision before smoothly entering the road, while PDM-C collides. On the right, we observe an agent about to cross the road. PDM-C drives slowly and collides, whereas our planner safely navigates forward. Please see our project page for videos.
@@
Oracle Target Parameters vs. BehaviorNet Predicted Parameters. We assess the quality of BehaviorNet’s predictions by comparing them to target IDM parameters from Eq. 1. The planning performance of AdaptiveDriver w/ Log-BehaviorNet closely approximates that of BehaviorNet w/ Log-Oracle, with differences of $0.14$ and $0.12$ in NR-CLS and R-CLS, respectively. The minimal difference between AdaptiveDriver using City-Oracle vs. City-BehaviorNet, and Log-Oracle vs. Log-BehaviorNet in Table IV, suggests that our approach can reliably predict future agent actions.
@@
AdaptiveDriver Generalizes to New Cities. As illustrated in Figure 4 (b), we observe that many agent behaviors are shared across cities. To validate this observation, we train AdaptiveDriver w/ Log-BehaviorNet on three cities and test it on a held-out city (cf. Table V). We find that the planning scores on NR-CLS and R-CLS for seen-vs-unseen cities are comparable, suggesting that AdaptiveDriver can generalize beyond city boundaries.
V Conclusion
@@
When does AdaptiveDriver Fail? Since our world models build on IDM’s PID controller, we inherit its flaws. We note that IDM can be too conservative (when it mistakes a parked vehicle for a lead vehicle and stops) or too aggressive (when traveling at high speeds along a curve. Secondly, IDM only interacts with the lead vehicle, limiting multi-agent reasoning. Although we primarily focus on building city-specific and log-specific world models using the IDM, modeling agent-specific behavior models can potentially yield better performance. Future work should explore alternative optimization functions to maximize distribution similarity between real driving logs and simulated rollouts for learning control parameters. We include additional videos on our project page.
Why Directly Predict Control Parameters? Directly predicting control parameters allows our world model to more accurately simulate multi-agent interactions in rollouts, improving proposal scoring. Further, relying on a motion controller allows us to more easily capture nuanced behaviors that may be difficult to learn like speeding up when merging in rollouts. However, relying on a rule-based controller for rollouts also implies that we inherit its limitations. Moreover, unlike learning-based approaches that directly predict agent behaviors, our approach may not significantly improve with more data. Given enough data and the right learning objectives, we posit that methods that directly predict agent behaviors may be able to mimic the characteristics of MPC-based controllers.
BehaviorNet Rollout Accuracy vs. Runtime BehaviorNet is limited by IDM’s lateral capability, preventing it from simulating multi-agent interactions like lane changes and lane merges. In theory, one could use more sophisticated controllers that allow for lang changes and merges that differ from the reference path (like the controller in PDM-C). Alternatively, one could replace a motion controller with an ego-forecaster such as MTR++ [42] or QCNet [43]. We believe both directions are fruitful paths forward for building better world models. In practice, we find that a simple IDM-based world model already achieves state-of-the-art performance. Moreover, similar to PDM-C’s constant velocity world model, our rule-based rollouts allow AdaptiveDriver to operate with lower latency. In contrast, running a SOTA trajectory prediction model at every timestep will significantly increase the latency of proposal scoring and is likely prohibitively slow for practical applications.
Learning-Based Rollouts Have Higher Latency We benchmark the latency of PDM-C and AdaptiveDriver on 1 RTX 3090 GPU with a batch size of 1 in Table VI. Notably, BehaviorNet only increases AdaptiveDriver’s runtime by 17ms compared to PDM-C.
                Model                 Latency (ms)                 PDM-C                 232                 AdaptiveDriver w/ Log-BehaviorNet                 249
TABLE VI: Latency Analysis. We compare the latency of PDM-C and AdaptiveDriver w/ Log-BehaviorNet. Unsurprisingly, AdaptiveDriver is slower than PDM-C due to our learning-based rollouts.
VI Acknowledgements
This work was supported in part by funding from Bosch Research and the NSF GRFP (Grant No. DGE2140739).